If the radius of a space station is 50m, what angular velocity is necessary so that a person of mass 80kg is subjected to a centrifugal force against the floor of the space station which equals their weight force on earth?
Answer #1

We need angular velocity of centrifuge for which centripetal force is equal to weight of person, So

Fc = W

m*w^2*R = m*g

w^2 = g/R

R = radius of space station = 50 m

w = sqrt (g/R) = sqrt (9.81/50)

w = 0.44 rad/sec = angular velocity required

Please Upvote.